What is the probability density of a continuous uniform distribution with boundaries at 3 and 8?

a

1/2

b

1/3

c

1/4

d

1/5
Answer #1

X ~ Uniform ( a , b)

Where a = 3 , b = 8

The probability density for uniform distribution = 1 / ( b - a)

= 1 / ( 8 - 3)

= 1 / 5